Math 125I Winter 2003 Name:

Quiz Section:

Quiz #3

SHOW YOUR WORK

1. (5 points) The graph of the function f(x) given in the figure consists of line segments and a semicircle. Let g(x) = ∫_{0}^{x} f(u) du.

(a) Compute g(2) and g(5)

(b) What is the maximum value of g(x) on the interval (0, 6)?

2. (5 points) Compute the definite integral ∫_{2}^{e} \frac{1}{x \sqrt{ln x}} dx